OpenStudy (anonymous):

Simplify the expression csc^2x sec^2x/ sec^2 x + csc^2 x

OpenStudy (anonymous):

This?\[ \large \frac{\csc^2(x) \sec^2(x)}{\sec^2(x)} + \csc^2(x) \]

OpenStudy (anonymous):

csc^2 is in the denominator with sec^2

OpenStudy (sirm3d):

\[\Large \frac{ \csc^2x \sec^2x }{ \sec^2x + \csc^2 x }\]

OpenStudy (anonymous):

Yes thats it

OpenStudy (sirm3d):

what basic identity can you remember that may help you simplify this problem?

OpenStudy (anonymous):

reciprocal identity

OpenStudy (sirm3d):

sub them in the problem.

OpenStudy (anonymous):

Im not really sure how to, can you please explain it?

OpenStudy (anonymous):

\[ \sec^2(x) = \frac{1}{\cos^2(x)} \]So how about multiplying top and bottom bye \(\cos^2(x)\).

OpenStudy (sirm3d):

\[\Huge \frac{ \frac{ 1 }{ \sin^2x }\frac{ 1 }{ \cos^2x } }{ \frac{ 1 }{ \cos^2 x }+\frac{ 1 }{ \sin^2x } }\]

OpenStudy (sirm3d):

if you are familiar with @wio 's method, you can multiply both sides of the fraction by \[\Large \sin^2x \cos^2x\]

OpenStudy (anonymous):

I am not can you explain it?

OpenStudy (sirm3d):

\[\frac{ \csc^2x \sec^2x }{ \sec^2x + \csc^2x }\cdot \frac{ \sin^2x \cos^2x }{ \sin^2x \cos^2 x }\]

OpenStudy (sirm3d):

multiply the numerators, multiply the denominators